A sample of gas contains 0.1900 mol of H2(g) and 9.500×10-2 mol of O2(g) and occupies a volume of 18.0 L. The following reaction takes place: 2H2(g) + O2(g)=2H2O(g) Calculate the volume of the sample after the reaction takes place, assuming that the temperature and the pressure remain constant.
Answer #1

Balanced reaction is

2H2 + O2  \rightarrow 2H2O

According to balanced reaction 2 mole H2 react with 1 mole O2 molar ratio between H2 to O2 is 2:1 thus to react with 0.1900 mole H2 required mole of O2 = 0.1900 / 2 = 9.5 X 10-2 mole so reactant given in proportion.

2 mole H2 produce 2 mole H2O molar ratio between H2 to H2O is 1:1 therefore 0.1900 mole H2 produce 0.1900 mole H2O

mole of product H2O produced = 0.1900 mole

Total mole of reactant = (mole of H2) + (mole of O2) = (0.1900) + (9.500 X 10-2) = 0.285 mole

Total mole of reactant = 0.285 mole

We know the Avogadro's law at constant temperature and pressure volume is directly proportional to no. of moles

V1/n1 = V2/n2

Where,

V1 = initial volume = 18 L

n1 = initial mole = 0.285 mole

V2 = final volume =?

n2 = final mole = 0.1900 moles

We can write above formula as

V2 = V1n2/ n1

Substitute the value

V2 = 18 L X 0.1900 mol / 0.285 mol = 12 L

After reaction take place volume of gas = 12 L